OpenStudy (anonymous):

Help Please! *METAL Evaluate the expression: v ⋅ w Given the vectors: r = <5, -5, -2>; v = <2, -8, -8>; w = <-2, 6, -5>

OpenStudy (anonymous):

The dot product of two vector u = <a, b, c> and k = <d, e, f> is \(u \cdot k = ad + be + cf\)

OpenStudy (anonymous):

As you can see, the dot product return a scalar. While a cross product will return a vector

OpenStudy (anonymous):

oh ok.

OpenStudy (anonymous):

2(-2) + (-8)6 + (-8)-5

OpenStudy (anonymous):

= -12 Right?

OpenStudy (anonymous):

So?

OpenStudy (anonymous):

That's correct!

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Np ;)

OpenStudy (anonymous):

Do you have time for another prob.

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

A certain radioactive isotope has a half-life of 15 days. If one is to make a table showing the half-life decay of a sample of this isotope from 32 grams to 1 gram; list the time (in days, starting with t = 0) in the first column and the mass remaining (in grams) in the second column, which type of sequence is used in the first column and which type of sequence is used in the second column?

OpenStudy (anonymous):

omg, my computer... ugh.

OpenStudy (anonymous):

The single element radioactive decay law is: \(dN/dt = - \lambda N\), where \(N\) is the mass at time t and \(\lambda\) is the decay constant of the element

OpenStudy (anonymous):

That's a first order ordinary differential equation of simple solution, just integrate: \(N(t) = N_0 e^{- \lambda t} \)

OpenStudy (anonymous):

Where \(N_0\) is teh initial mass

OpenStudy (anonymous):

Well, let's find a ralation of \(\lambda\) with the half-life \(t_{1/2}\)

OpenStudy (anonymous):

\(N_0 /2 = N_0 e^{- \lambda t}\), now find t that represents the half-life: \[t = t_{1/2} = \frac{ln(2)}{\lambda}\]

OpenStudy (anonymous):

Now, substitute it back on the equation for N. \[N(t) = N_0 e^{- \frac{ln(2) t}{t_{1/2}}}\]

OpenStudy (anonymous):

Now, with this equation, you can calculate the mass of the element for every time t.

OpenStudy (anonymous):

in order to make easier to make the table, set \(t = k*t_{1/2}\), where k = 0, 1, 2, ...